OpenStudy (abbycross167):

Write the point-slope (y - k = m(x - h)) form of the equation from the information given. 6. Slope = -½; (h, k) = (-1, 2) Can someone please show me how to answer this?

pooja195 (pooja195):

y-k=m(x-h) m=1/2 h=-1 k=2

OpenStudy (abbycross167):

oh wow lol that was fast... thank you very much, could help me with 2 more?

pooja195 (pooja195):

Sure =)

OpenStudy (abbycross167):

Slope equals 2 and the line goes through the point (1, -3)

OpenStudy (abbycross167):

i don't understand how to do this one

pooja195 (pooja195):

In point slope form?

OpenStudy (abbycross167):

umm i guess so :/

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (abbycross167):

is that all?

pooja195 (pooja195):

es Y

pooja195 (pooja195):

*yes

OpenStudy (abbycross167):

Wouldn't I have to add some of the numbers?

pooja195 (pooja195):

Does the question ask what form they want it in?

pooja195 (pooja195):

I assume point slope form because thats what they asked for in the first one.

OpenStudy (abbycross167):

I need to add the numbers...

OpenStudy (abbycross167):

@pooja195

OpenStudy (anonymous):

y-2=-1/2(x+1)

OpenStudy (abbycross167):

Thank you!! @TheSavior